I almost have this. I just have a problem filling in the formulas from word problems. the word problem reads: Martina leaves home at 9am bicycling at a rate of 24mph. 2 hours later, john leaves driving at the rate of 48 mph. at what time will john catch up with martina. i had 48=24(2x). the equation is true.. i dont if it is right for this problem.
Answer by Paul(988)

Let the time be x
24(x+2)=48x
24x+48=48x
24x=28
x=2
9:00 +2 = 11:00
Hence, John will catch Matrina at 11:00am.
PAUL.